OpenStudy (anonymous):

Emilio's age is 4 yrs old more than half the age of his sister Rosa. The sum of their ages is 28. How old is Emilio?

OpenStudy (jiteshmeghwal9):

Let the age of emilo's sister be 'x'.then Emilo's age=x+4 nw, A/q \[(x)+(x+4)=28\]\[2x+4=28\]\[2x=28-4\]\[2x=24\]\[x=12\]